OpenStudy (anonymous):

what is f(x+h)-f(x) when f(x)=3x^2+x-2

OpenStudy (anonymous):

@ganeshie8

OpenStudy (butterflydreamer):

f (x+h) just means that for every "x" in f (x) = 3x^2 + x - 2 You would want to repalce the "x" with (x + h) f (x) = 3x^2 + x - 2 Therefore, f (x + h) = 3(x+h)^2 + (x+h) - 2

OpenStudy (anonymous):